### v2

# Connecting Star Formation in the Milky Way and Nearby Galaxies -II. An Observationally Driven Analytical Model for Predicting Cloud-Scale Star Formation Rates

Abstract
We construct a model by integrating observational constraints from the Milky Way and nearby galaxies to predict cloud-scale star formation rates (SFRs). In the model, we first estimate the initial total mass of clumps in a cloud based on the cloud mass, and then generate the initial clump population of the cloud using the initial clump mass function. Next, we model the star formation histories (SFHs) of the cloud to assign an age to each clump. We then sort out the intermediate-age clumps and calculate the total embedded cluster mass. Finally, we predict the SFR based on the duration of the embedded phase. The model-predicted SFR is broadly consistent with the observed SFR, supporting the plausibility of the model. The model primarily provides a theoretical framework that integrates a wide range of observational results, thereby clarifying the tasks for future observations.
keywords:
– ISM: clouds – galaxies: clusters – galaxies: star formation
pubyear: 2025
1 Introduction
Galaxies act as stellar nurseries, forming stars through the gravitational collapse of dense regions within molecular clouds. These clouds, which are widespread throughout galaxies, contain multiple clumps that serve as the local sites of star formation and the precursors of embedded star clusters (Kennicutt & Evans, 2012; Miville-Deschênes et al., 2017; Rosolowsky et al., 2021; Urquhart et al., 2022; Yan et al., 2017; Zhou et al., 2024d, c). The work of Motte et al. (2018); Vázquez-Semadeni et al. (2019); Kumar et al. (2020); Henshaw et al. (2020); Zhou et al. (2022, 2023, 2024b); Zhou & Davis (2024); Zhou et al. (2024a); Zhou & Li (2025); Zhou et al. (2025b) (and reference therein) present a comprehensive multi-scale investigation of hub-filament structures, examining their morphology, kinematics, and evolution from dense core ($\sim$1000 AU) to clump ($\sim$1 pc), molecular cloud ($\sim$10-100 pc), and galaxy-cloud ($\sim$1000 pc) scales. These observations reveal hierarchical hub-filament structures with self-similar properties spanning sub-parsec to kiloparsec scales, highlighting their crucial role in the star formation process. Within this hierarchy, a dense core serves as a hub within a clump, a clump functions as a hub within a molecular cloud, and a molecular cloud acts as a hub within a galaxy. Hub-filament structures form through the gravitational contraction of gas structures. Velocity gradient analyses suggest that gas inflows along filaments are driven by gravity. Molecular gas is organized into network structures formed by the gravitational coupling of multi-scale hub-filament structures, where local gravitational centers (hubs) serve as the primary sites of star formation.
Progress in understanding the physical properties of giant molecular clouds (GMCs) is critical for elucidating the connection between interstellar gas dynamics and star formation on galactic scales. Understanding what regulates the star formation rate (SFR) in molecular clouds is a central problem in star formation theory. Despite the short gravitational free-fall times of dense gas, observations show that star formation is remarkably inefficient, with only a few percent of the gas converted into stars (Evans et al., 2009; Kennicutt & Evans, 2012; Kim et al., 2022; Chevance et al., 2023; Zhou et al., 2025a). A wide variety of theoretical models have been proposed to explain this inefficiency, invoking turbulence, magnetic fields, global gravitational collapse, and stellar feedback. In turbulence-regulated models, supersonic (magneto-)hydrodynamic turbulence generates a broad density distribution, and only gas in the high-density tail becomes gravitationally unstable (Krumholz & McKee, 2005; Padoan & Nordlund, 2011; Hennebelle & Chabrier, 2011; Federrath & Klessen, 2012). Alternative models emphasize the time-dependent nature of molecular clouds, proposing that star formation proceeds via global, hierarchical collapse and accelerates as clouds evolve (Zamora-Avilés & Vázquez-Semadeni, 2014; Vázquez-Semadeni et al., 2019). Observationally motivated approaches link the star formation rate directly to the mass of dense gas ($M_{\rm dense}$) above a column density threshold, implying $\mathrm{SFR}\propto M_{\rm dense}$ (Lada et al., 2010; Shimajiri et al., 2017). Feedback-regulated models argue that stellar feedback injects momentum and energy that balance gravity on cloud scales, self-adjusting the star formation rate, particularly in massive or high-surface-density clouds (Ostriker & Shetty, 2011; Kruijssen et al., 2019).
Recent advances in high-resolution, multi-wavelength observations now allow star formation to be resolved on the scale of individual molecular clouds in nearby galaxies. In particular, high-resolution CO imaging from ALMA (Atacama Large Millimeter/submillimeter Array), along with other submillimeter observations, enables systematic investigations of the molecular cloud population beyond the Milky Way (Leroy et al., 2021a, b; Lee et al., 2022; Emsellem et al., 2022; Lee et al., 2023; Grishunin et al., 2024; Schinnerer & Leroy, 2024). Since molecular clouds in the Milky Way can be resolved down to their internal structures, they provide a crucial reference for understanding the internal composition and star formation processes of molecular clouds in nearby galaxies. Star formation occurs primarily within clumps. Star formation observed on molecular cloud scales or larger is essentially an integrated outcome of the activity occurring within these clumps. To truly understand the underlying physics of star formation at those larger scales, it is essential to characterize the star-forming states of individual clumps. This requires turning to the Milky Way as a reference, leveraging surveys such as ATLASGAL and Hi-GAL (Schuller et al., 2009; Urquhart et al., 2022; Molinari et al., 2010; Elia et al., 2021), which are capable of resolving individual clumps, along with numerous ALMA follow-up studies that zoom in on single clumps in detail (Sanhueza et al., 2019; Liu et al., 2020; Motte et al., 2022; Molinari et al., 2025). In Zhou & Dib (2025), we used CO (2–1) and CO (1–0) data cubes to identify molecular clouds and study their kinematics and dynamics in three nearby galaxies and the inner Milky Way. Molecular clouds in the same mass range across these galaxies show broadly comparable physical properties and similar star formation rates. Strong correlations were found between cloud mass and total clump mass, clump number, and the mass of the most massive clump (see also Zetterlund et al., 2019).
Unlike previous theoretical models, this work adopts a predominantly observationally driven approach. Rather than relying on specific theoretical assumptions, we attempt to synthesize existing observational results and treat the molecular-cloud-scale star formation rate as a natural consequence of a set of empirical relations. Clumps are identified as the true sites of star formation within molecular clouds. The star formation rate on molecular cloud scales, as a large-scale phenomenon, is in fact an integrated manifestation of star formation processes occurring on clump scales within molecular clouds. As self-gravitating structures (Liu et al., 2016; Urquhart et al., 2018; Evans et al., 2021), clumps are almost decoupled from their surrounding large-scale environments (Watkins et al., 2019; Peretto et al., 2023; Zhou et al., 2023, 2024f) and therefore form stars in a nearly independent manner. Consequently, clumps, rather than molecular clouds, are taken to be the fundamental units of star formation. The structure of this paper is as follows: In Section 2, we compile empirical relations at the molecular cloud and clump scales derived from observations of the Milky Way and nearby galaxies, construct the initial clump mass function of molecular clouds, and model their star formation histories. Section 3 describes in detail the procedure for implementing the model predictions, presents an uncertainty analysis of the model parameters, and discusses the limitations of the model. Section 4 provides a brief summary.
2 Model
Table 1: Notation and symbol definitions used in the model.
Symbol Description
$M_{\rm cloud}$ Total mass of a molecular cloud
$M_{\rm clump}$ Mass of an individual clump within a cloud
$M_{\rm clump,tot}$ Initial total mass of clumps in a cloud (including past and present clumps)
$M_{\rm clump,tot,obs}$ Observed total mass of clumps in a cloud (currently existing clumps only)
$M_{\rm clump,tot,m}$ Total mass of clumps in the transition model between low- and high-mass clouds (Equation 14)
$M_{\rm ecl}$ Mass of an embedded star cluster
$\mathrm{SFR}_{\rm cloud,obs}$ Observed star formation rate of a molecular cloud
$\mathrm{SFR}_{\rm cloud,p}$ Model-predicted star formation rate of a cloud
$\Sigma_{\rm SFR}$ Star formation rate surface density (SFR per unit area)
[…]
AA Projected area of a molecular cloud
$\mathrm{SFE}_{\rm clump}$ Star formation efficiency of a clump
$\mathrm{SFE}_{\rm cloud}$ Star formation efficiency of a cloud (stellar mass / cloud mass)
$\xi_{\rm clump}(M)$ Clump mass function (CLMF)
$\beta$ Power-law index of the clump mass function
$M_{\rm min}$ Minimum clump mass
$M_{\rm max}$ Maximum clump mass
$t_{\rm b}$ Age of a clump
$t_{\rm b,p}$ Peak of the clump age distribution (Gaussian SFH)
$\sigma_{\rm t}$ Standard deviation of clump age distribution (Gaussian SFH)
$t_{\rm GMC}$ Typical lifetime of a giant molecular cloud (GMC)
$t_{\rm emb}$ Duration of the embedded phase of a cluster
$\overline{\rm CV}_{i}$ Average coefficient of variation for parameter $i$ (measure of relative uncertainty)
$\overline{\Delta}_{i}$ Average relative deviation of parameter $i$ from the reference SFR
$f_{i}$ Fractional contribution of parameter $i$ to total uncertainty
$k_{\rm clump}$ Normalization constant of the CLMF
2.1 Correlations from the Milky Way
Figure 1: Correlations between the total clump mass in a cloud ($M_{\mathrm{clump,tot,obs}}$), the cloud mass ($M_{\mathrm{cloud}}$) and the SFR of the cloud ($\mathrm{SFR}_{\mathrm{cloud,obs}}$) in the inner Milky Way. The red dashed line represents a linear fit to the scatter points, with $k$ and $r$ denoting the slope and the correlation coefficient. The contours show the density distribution of the scatter points, derived from a Kernel Density Estimate (KDE).
[…]
$\frac{\Sigma_{\rm SFR}}{M_{\odot}~yr^{-1}~kpc^{-2}}=3.8\times 10^{-3}\left(\frac{I_{\rm 24\mu m}}{MJy~sr^{-1}}\right),$ (1)
$\mathrm{SFR}=\Sigma_{\rm SFR}*A$ (2)
[…]
$\begin{split}&\mathrm{log}_{\rm 10}(M_{\mathrm{clump,tot,obs}}/M_{\odot})=\\ &(0.75\pm 0.02)\times\mathrm{log}_{\rm 10}(M_{\mathrm{cloud}}/M_{\odot})-(0.56\pm 0.13),\end{split}$ (3)
$\begin{split}&\mathrm{log}_{\rm 10}(\mathrm{SFR}_{\mathrm{cloud,obs}}/(M_{\odot}~yr^{-1}))=\\ &(0.98\pm 0.01)\times\mathrm{log}_{\rm 10}(M_{\mathrm{cloud}}/M_{\odot})-(9.12\pm 0.06).\end{split}$ (4)
In Zhou et al. (2024c), for the ATLASGAL clumps, we found that the clump star formation efficiency ($\mathrm{SFE}_{\mathrm{clump}}$) decreases with increasing clump mass ($M_{\mathrm{clump}}$), with a median value of $\approx$0.3. We also derived the correlations between the $\mathrm{SFE_{\mathrm{clump}}}$, the clump mass and the embedded cluster mass ($M_{\mathrm{ecl}}$),
$\begin{split}&\mathrm{log}_{10}(\mathrm{SFE_{\mathrm{clump}}})=\\ &(-0.37\pm 0.01)\times\mathrm{log}_{10}(M_{\mathrm{clump}}/M_{\odot})+(0.42\pm 0.04),\end{split}$ (5)
$\begin{split}&\mathrm{log}_{10}(M_{\mathrm{clump}}/M_{\odot})=\\ &(1.02\pm 0.02)\times\mathrm{log}_{10}(M_{\mathrm{ecl}}/M_{\odot})+(0.52\pm 0.05).\end{split}$ (6)
2.2 Clump mass function
Given that clumps represent the progenitors of embedded star clusters, we assume a direct inheritance relationship between the clump mass function (CLMF) and the embedded cluster mass function (ECMF). Then we adopt the formalism summarized in Yan et al. (2017). The CLMF is a single slope power law with a variable power-law index, $\beta$,
$\xi_{\mathrm{clump}}(M)=k_{\mathrm{clump}}\times M^{-\beta},M_{\mathrm{min}}\leqslant M<M_{\mathrm{max}},$ (7)
where $M_{\mathrm{min}}$ is the assumed lower limit of clump masses, $M_{\mathrm{max}}$ is the upper integration limit in the optimal sampling method defined in Schulz et al. (2015) , and $k_{\mathrm{clump}}$ is a normalization constant. The mass of the smallest stellar group observed in Kuhn et al. (2015); Zhou et al. (2024e) is $\approx 5M_{\odot}$. When the uncertainties in both equations 5 and 6 are taken into account, the resulting SFE lies in the range [0.83, 1]. In the limiting case of SFE=1, the minimum clump mass is also 5 $M_{\odot}$. In order to guarantee sufficient sampling, we set the minimum clump mass to $M_{\mathrm{min}}=5M_{\odot}$. Generally, the slope of the star cluster mass function is $\approx$ -2 (Lada & Lada, 2003; Krumholz et al., 2019; Mok et al., 2020; Wainer et al., 2022; Zhou et al., 2024e, 2025a). In Elia et al. (2017), for Hi-GAL (Herschel InfraRed Galactic Plane Survey) clumps in the inner Milky Way, the slopes of the clump mass functions are between -1.88 and -2.46.
The parameters $k_{\mathrm{clump}}$ and $M_{\mathrm{max}}$ in equation 7 are determined by solving equations 8 and 9 together, i.e., by invoking the clump population mass conservation:
$M_{\mathrm{clump,tot}}=\int_{M_{\mathrm{min}}}^{M_{\mathrm{max}}}M~\xi_{\mathrm{clump}}(M)\,\mathrm{d}M.$ (8)
The optimal sampling normalization condition:
$1=\int_{M_{\mathrm{max}}}^{10^{9}~\mathrm{M}_{\odot}}\xi_{\mathrm{clump}}(M)\,\mathrm{d}M,$ (9)
see Yan et al. (2017) for details on the adopted upper integration limit of $10^{9}$ $M_{\odot}$. Actually, as long as the upper integration limit is much larger than the real $M_{\mathrm{max}}$, the result will not be affected. In Section 2.1, the mass of the most massive clump is only $10^{4.6}$ $M_{\odot}$.
2.3 Initial total mass of clumps
Clumps, which are self-gravitating structures within molecular clouds (Liu et al., 2016; Urquhart et al., 2018; Evans et al., 2021), evolve more rapidly than the clouds themselves. By the time large-scale molecular clouds condense into dense, localized regions within galaxies, star formation has often already taken place within these regions. The molecular clouds we observe are typically composites, containing both exposed and embedded stellar populations, as well as clumps (Kuhn et al., 2020; Turner et al., 2022; Zhou et al., 2025a). Nearly all of those stellar populations—whether visible or still embedded—originate from clumps. These clumps may either be currently existing or have existed in the past. Past clumps are now observed as exposed and embedded stellar populations. Together, existing and past clumps constitute the initial clump population of the cloud, described by the initial clump mass function introduced in Section 2.2. Therefore, $M_{\mathrm{clump,tot}}$ in equation 8 and $M_{\mathrm{clump,tot,obs}}$ in equation 3 are different. $M_{\mathrm{clump,tot}}$ includes both past and present clumps, while $M_{\mathrm{clump,tot,obs}}$ includes only the clumps currently observed. As an estimate,
$M_{\mathrm{clump,tot}}$ $=M_{\mathrm{clump,tot,obs}}+(M_{\mathrm{emb,tot,obs}}+M_{\mathrm{exp,tot,obs}})/0.3$
$=M_{\mathrm{clump,tot,obs}}+M_{\mathrm{cloud}}\times\mathrm{SFE}_{\mathrm{cloud}}/0.3,$ (10)
[…]
$\begin{split}&\mathrm{log}_{\rm 10}(\mathrm{SFE}_{\mathrm{cloud}})=\\ &(-0.32\pm 0.06)\times\mathrm{log}_{\rm 10}(M_{\mathrm{cloud}}/M_{\odot})-(0.034\pm 0.37).\end{split}$ (11)
2.4 Star formation history
Equation 7 defines the initial clump mass function of the cloud. The theoretically initial clump population has now evolved into a mixture of the currently observed clumps and the embedded/exposed stellar populations. The currently observed clumps correspond to the youngest ones, while the embedded and exposed stellar populations can be regarded as the intermediate-age and oldest clumps, respectively. The different components in a molecular cloud — clumps, and embedded and exposed stellar populations — can be regarded as representing the age spread of clumps within the cloud. We then model the star formation history (SFH) of the cloud to distinguish between these different components. Using the SFH models presented in Zhou et al. (2024d); Dib et al. (2025), we examine scenarios where the SFH remains constant and others where it varies over time. Through the SFH, we create an age distribution of the initial clump population in a cloud to assign an age to each clump. Then, we sort out the intermediate-age clumps, which represent the embedded stellar populations and mainly contribute to the mid-infrared emission of the cloud.
The typical survival timescale of GMCs ($t_{\rm GMC}$) is 10-30 Myr (Chevance, 2021; Kim et al., 2021, 2022). In the cases of a constant SFH, we randomly sample the age of clumps, $t_{\rm b}$, with a uniform probability in the age range [0, $t_{\rm GMC}$] Myr. For a time-varying SFH, we employ a Gaussian function, allowing us to adjust both the peak position and the distribution width in time. The clump age, $t_{\rm b}$, is then randomly drawn from the corresponding Gaussian distribution,
$P(t_{\rm b})=\frac{1}{\sigma_{\rm t}\sqrt{2\pi}}\exp\left[-\frac{1}{2}\left(\frac{t_{\rm b}-t_{\rm b,p}}{\sigma_{\rm t}}\right)^{2}\right],$ (12)
where $t_{\rm b,p}$ is the peak position, and $\sigma_{\rm t}$ is the standard deviation. We should consider different values of $\sigma_{\rm t}$ to cover possible time spans of star formation. We consider values of $t_{\rm b,p}=2,4,6,8$ Myr and $\sigma_{\rm t}=$ 3, 6 and 9 Myr. There are a total of 12 cases here. Actually, the time-dependent SFH with large $\sigma_{\rm t}$ is similar to the constant SFH. From a clump to an embedded cluster, the formation time is $\approx$2 Myr (Evans et al., 2009; Covey et al., 2010; Megeath et al., 2022; Wells et al., 2022; Kim et al., 2023). Thereafter, the duration time of the embedded phase ($t_{\rm emb}$) is $\approx$2-7 Myr (Kim et al., 2021, 2023). Therefore, we only select the clumps in the age range [2, $t_{\rm emb}$+2] Myr to calculate the SFR and compare with the observation. Equation 6 is used to convert the clump mass to the embedded cluster mass. Then the total embedded cluster mass is divided by $t_{\rm emb}$ to predict the SFR of the cloud ($\mathrm{SFR}_{\mathrm{cloud,p}}$).
3 Results and discussion
3.1 Procedure
The predictions of the model account for the cumulative effect of uncertainties across the entire modeling chain. Rather than providing single-point estimates, the model generates probability distributions for SFR predictions, acknowledging that our knowledge of physical parameters is inherently uncertain. The model adopts a nested Monte Carlo framework in which molecular cloud masses are sampled over a logarithmic grid from $10^{3}$ to $10^{8}$ M (1000 points). For each cloud mass, multiple Monte Carlo realizations (typically 100) are performed, with all model parameters independently resampled in each iteration. This design enables a systematic assessment of how uncertainties propagate across the cloud mass spectrum. The resulting ensemble of SFR predictions for each mass is used to derive statistical descriptors, including the mean and median as central estimates and the standard deviation as a measure of dispersion. A key strength of this approach is the full propagation of uncertainties through all stages of the calculation, naturally accounting for non-linear effects and parameter correlations that are not captured by simplified analytic error-propagation methods.
The logic and procedures of the model are as follows: (1) Use equation 10 to estimate the initial total mass of clumps in a cloud; (2) Generate the initial clump population of the cloud using the initial clump mass function; (3) Model the SFH of the cloud to assign an age to each clump; (4) Identify the clumps in the age range [2, $t_{\rm emb}$+2] Myr and calculate the total embedded cluster mass; (5) Predict the SFR based on the duration of the embedded phase. Some details should be emphasized here: (1) We sample $\beta$ uniformly between 1.88 and 2.46 in each Monte Carlo iteration. This reflects the observed variation in CLMF slopes across different molecular clouds and star-forming regions. The uniform distribution represents a conservative assumption of maximum uncertainty within the observed range; (2) The maximum age for star formation history is sampled uniformly between 10 and 30 Myr, reflecting the typical survival timescale of GMCs. The starting point of the embedded phase is fixed at 2 Myr, representing the time required for a clump to evolve into an observable embedded cluster. The endpoint of the embedded phase is sampled uniformly between 4 and 9 Myr, since the duration time of the embedded phase is $\approx$2-7 Myr; (3) All empirical relations derived from observations in Section 2 are parameterized with truncated normal priors on their slopes and intercepts, capturing empirical uncertainties while enforcing physically meaningful bounds.
The model incorporates multiple internal consistency checks. Mass conservation verifies that the generated clump masses sum to the specified total mass within tolerance. Physical plausibility checks ensure that converted embedded cluster masses never exceed their parent clump masses. Age distribution checks confirm that assigned ages respect the specified SFH model. For each molecular cloud mass, we monitor convergence of the Monte Carlo statistics as iterations progress. We check stabilization of the mean, standard deviation, and percentile estimates to ensure sufficient sampling of the parameter space. The model reports the fraction of successful iterations (those completing without numerical errors) as a quality metric, see the online implementation codes for more details  [^1] https://github.com/jianwenzhou11/SFR_model.git.
3.2 Transition mass
Figure 2: Compare the observed SFR with the model-predicted SFR. Green points represent the observation, the black dashed line shows the linear fit to the green points, with $k$ and $r$ denoting the slope and the correlation coefficient. (a) The orange and blue points are the predictions under the constant SFH based on equation 10 and equation 3, respectively, and the shaded regions indicate the uncertainty ranges; (b) Solid and dashed lines indicate the linear fits of the predictions for the 13 different SFHs, with the red line highlighting the result for the constant SFH. The black dashed line represents the observation; (c) Same as panel (a). The red dashed line shows a fit to the predictions based on equation 14, while the cyan shaded region indicates the uncertainty in the predicted results.
Fig.2(a) shows the results under a constant SFH. At the high-mass end, $\mathrm{SFR}_{\mathrm{cloud,p}}$ and $\mathrm{SFR}_{\mathrm{cloud,obs}}$ are comparable, but $\mathrm{SFR}_{\mathrm{cloud,p}}$ is significantly larger than $\mathrm{SFR}_{\mathrm{cloud,obs}}$ at the low-mass end. If we change equation 10 into equation 3, the result is opposite. All 13 SFHs (12 Gaussian and 1 constant) are compared in Fig.2(b), and the constant SFH simultaneously provides the best fit to both the low-mass and high-mass ends. Therefore, we focus exclusively on the constant SFH in the following discussion.
The discussion in Section 2.3 mainly concerns molecular clouds at the high-mass end, which can survive for a long time despite hosting stellar populations. In contrast, in low-mass molecular clouds, feedback from the stellar populations can efficiently disrupt the cloud. Hence, observed low-mass molecular clouds are not expected to contain a significant fraction of stellar components. Actually, equation 11 from Zhou et al. (2025a) is only for clouds with masses larger than $10^{5.5}$ $M_{\odot}$. Therefore, we need a new relation: it should follow equation 10 for high-mass molecular clouds, equation 3 for low-mass molecular clouds, and lie between the predictions of the two equations for intermediate-mass molecular clouds. To obtain a smooth transition between the two regimes, we adopt a linear weighting function defined as
$t=\frac{\mathrm{log}_{\rm 10}(M_{\mathrm{cloud}}/M_{\odot})-a}{b-a},$ (13)
where $a$ and $b$ specify the transition mass range. $t$ is limited to the range [0,1]. The final relation is then given by
$M_{\mathrm{clump,tot,m}}=(1-t)\,M_{\mathrm{clump,tot,obs}}+t\,M_{\mathrm{clump,tot}},$ (14)
where $M_{\mathrm{clump,tot,obs}}$ and $M_{\mathrm{clump,tot}}$ correspond to the relations of equation 3 and equation 10, respectively. Using only equation 10 or equation 3 represents two extreme cases. A grid of molecular cloud masses spanning the full range of interest ($10^{3}$, $10^{8}$] M) is constructed, and for each mass, Monte Carlo simulations are performed assuming purely equation 3 or purely equation 10 contributions. By fitting the two extremes, we estimate the critical masses at which the model predictions begin to significantly deviate from the observation and then perform local optimization starting from these initial values. The optimization seeks to drive the median of the ratio between the predicted and target SFRs toward unity while minimizing its variance. The loss function combines an absolute-deviation term with a standard-deviation penalty:
$\mathcal{L}=|R_{\text{median}}-1|+\frac{\sigma_{R}}{R_{\text{median}}}$ (15)
where $R_{\text{median}}$ is the median SFR ratio across all Monte Carlo realizations, and $\sigma_{R}$ is the standard deviation of the SFR ratio. Finally, we obtain $a\approx 4$ and $b\approx 6.5$. As shown in Fig.2(c), after the combination of equation 10 and equation 3, now the model can well fit the observation.
3.3 Uncertainty
Table 2: Parameter-specific contributions to the total uncertainty ranked by $\overline{\text{CV}}_{i}$ in descending order.
Parameter $\overline{\text{CV}}_{i}$ $\overline{\Delta}_{i}$ $f_{i}$
equation 11 0.62 0.37 15.94
equation 11_slope 0.60 0.36 15.42
$t_{\rm GMC}$ 0.43 0.36 11.05
equation 3_slope 0.30 0.32 7.71
equation 6_intercept 0.30 0.28 7.71
$t_{\rm emb}$ 0.29 0.33 7.46
equation 6_slope 0.28 0.31 7.19
equation 4_intercept 0.27 0.28 6.94
$\beta$ 0.27 0.28 6.94
equation 3_intercept 0.27 0.30 6.94
equation 4_slope 0.26 0.26 6.68
The uncertainty analysis is based on a One-at-a-Time (OAT) methodology, with additional consideration of statistical sampling and parameter distributions. Before evaluating the contributions of individual parameters, a reference is defined using the median values of all parameters. Specifically, for the reference, we neglect uncertainties in all empirical relations, and for parameters defined within a finite range, we adopt the midpoint of the allowed interval. Each parameter $p_{i}$ is then examined independently, while all other parameters are held fixed at their reference values. This method can decompose the total uncertainty in SFR predictions into contributions from individual parameters and physical processes. It is crucial for identifying which sources of uncertainty dominate the model’s predictive variance and guiding efforts to improve model precision.
For each parameter $p_{i}$ and each cloud mass $M_{k}$, the set of SFR predictions $\{\text{SFR}_{i}^{(j)}(M_{k})\}_{j=1}^{N}$ is analyzed:
$\mu_{i}(M_{k})$ $=\frac{1}{N}\sum_{j=1}^{N}\text{SFR}_{i}^{(j)}(M_{k}),$
$\sigma_{i}(M_{k})$ $=\sqrt{\frac{1}{N-1}\sum_{j=1}^{N}\left(\text{SFR}_{i}^{(j)}(M_{k})-\mu_{i}(M_{k})\right)^{2}},$
$\text{CV}_{i}(M_{k})$ $=\frac{\sigma_{i}(M_{k})}{\mu_{i}(M_{k})}\quad\text{(for $\mu_{i}>0$)},$
$\Delta_{i}(M_{k})$ $=\frac{|\mu_{i}(M_{k})-\text{SFR}_{\text{ref}}(M_{k})|}{\text{SFR}_{\text{ref}}(M_{k})},$
where $i$ denotes the $i$-th parameter in the analysis. $j$ indexes the $j$-th random sample drawn for parameter $i$ at cloud mass $M_{k}$. $\text{SFR}_{i}^{(j)}(M_{k})$ is the predicted SFR obtained when parameter $i$ is set to its $j$-th randomly sampled value, while all other parameters are held fixed at their reference values. $N$ represents the Monte Carlo sample size. Then the statistics are integrated across the cloud mass range:
$\overline{\text{CV}}_{i}=\frac{1}{N_{\text{Mcloud}}}\sum_{k=1}^{N_{\text{Mcloud}}}\text{CV}_{i}(M_{k}),$ (16)
$\overline{\Delta}_{i}=\frac{1}{N_{\text{Mcloud}}}\sum_{k=1}^{N_{\text{Mcloud}}}\Delta_{i}(M_{k}).$ (17)
Parameter-specific contributions to the total uncertainty are then ranked by $\overline{\text{CV}}_{i}$ in descending order. $\overline{\text{CV}}_{i}$ is prioritized over $\overline{\Delta}_{i}$ because it directly quantifies the prediction uncertainty range—how much SFR predictions scatter due to parameter variations—whereas $\overline{\Delta}_{i}$ only measures systematic bias from a reference. In uncertainty analysis, understanding the full spread of possible outcomes (captured by $\overline{\text{CV}}_{i}$) is more critical than knowing the average offset, as symmetric distributions can show zero $\overline{\Delta}_{i}$ while still having substantial uncertainty. The scale-invariance of $\overline{\text{CV}}_{i}$ also allows consistent comparison across parameters despite the five-order magnitude range in $M_{\rm cloud}$, making it the appropriate metric for ranking parameter contributions to overall model uncertainty.
The final step decomposes total uncertainty into contributions from individual parameters, employing the coefficient of variation as a dimensionless and scale-invariant measure:
$f_{i}=\frac{\overline{\text{CV}}_{i}}{\sum_{j=1}^{N_{\text{p}}}\overline{\text{CV}}_{j}},$ (18)
where $N_{\text{p}}$ is the number of parameters. As reflected in Table.2, equation 11 contributes $\approx$31 % of the total uncertainty, due to the large scatter in the molecular cloud SFE as a function of cloud mass. In Zhou et al. (2025a), the Pearson correlation coefficient of this equation is only $\approx$0.27. The significant uncertainty introduced by equation 11 is also clearly reflected in Fig.2(a), where the orange-shaded uncertainty band is substantially wider than the blue-shaded band.
3.4 Current Limitations and Caveats
In the model, we considered three components within the molecular cloud, namely clumps, embedded and exposed stellar populations. The exposed stellar populations have already separated from the gas and therefore do not significantly affect the physical properties of the clumps. As discussed in Watkins et al. (2019); Zhou et al. (2024f, g), although the embedded stellar populations are spatially closely connected to the clumps, the feedback driven by the embedded stellar populations do not significantly influence the clumps’ physical properties. The reason lies in the hierarchical structure of the molecular gas, as described in Section 1. The network structure of molecular clouds (Zhou et al., 2025b) implies that the knots or local dense structures as local hubs or local gravitational centers are relatively independent of each other. Although some knots in the cloud evolve more rapidly, forming embedded stellar populations, due to their relative independence from the neighboring knots, early feedback from them does not significantly impact the physical properties of the neighboring knots (clumps).
In Urquhart et al. (2022), 5007 ATLASGAL clumps are classified into four evolutionary stages (quiescent, protostellar, young stellar objects and HII regions) and find similar numbers of clumps in each stage. The ATLASGAL clumps with HII regions (HII-clumps) are considered to be in the final stage of star formation and contain fully formed embedded star clusters (Wells et al., 2022; Zhou et al., 2024d, c). Assuming that HII-clumps account for 1/4 of the total number of clumps, and that the typical star formation efficiency of clumps is $\approx$0.3 (Zhou et al., 2024c, and references therein), the fraction of stellar mass contained in the clumps is 1/4*0.3 (7.5%), and can therefore be neglected.
The uniform distributions for some parameters ($\beta$, $t_{\rm GMC}$, and $t_{\rm emb}$) represent maximum uncertainty assumptions rather than informed priors. The single-slope power-law CLMF may oversimplify the true mass distribution of clumps. The fixed starting point of the embedded phase at 2 Myr neglects possible environmental variations. In Section 3.2, the total clump mass predicted from the cloud mass significantly affects the results. To accurately estimate the initial total clump mass within a molecular cloud, we need precise measurements of both the current total clump mass and the total stellar mass contained within the cloud, especially for low-mass clouds. Equation 14 currently represents a compromise, and the exact relationship still needs to be constrained by observations.
Anyway, this model primarily provides a theoretical framework that integrates a wide range of observational results, thereby clarifying the tasks for future observations. At a minimum, observations are needed to constrain a set of model parameters that exhibit significant uncertainties, as shown in Table.2.
4 Conclusion
In the model, we compile empirical relations at the molecular cloud and clump scales, derived from observations of the Milky Way and nearby galaxies. Based on these relations, we construct the initial clump mass function (CLMF) of molecular clouds and model their star formation histories (SFHs). We then employ a nested Monte Carlo framework to predict cloud-scale star formation rates, fully propagating uncertainties from all model parameters. Cloud masses are sampled logarithmically from $10^{3}$ to $10^{8}~M_{\odot}$, and for each mass, multiple realizations resample parameters independently, producing probability distributions of SFRs rather than single-point estimates. Statistical descriptors, including the mean, median, and standard deviation, are used to quantify the central predictions and associated uncertainties. Parameters such as the CLMF slope ($\beta$), GMC lifetime ($t_{\rm GMC}$), and embedded phase duration ($t_{\rm emb}$) are sampled from uniform distributions to account for observed variations, while empirical relations are assigned truncated normal priors to capture measurement uncertainties and enforce physically meaningful bounds.
Comparison with observations indicates that a constant SFH provides the best overall fit at both low-mass and high-mass ends. A smooth transition function (equation 14) interpolates between these regimes by combining predictions from equations 3 and 10 (two extreme cases), with local optimization identifying the transition mass range, $\log_{10}(M_{\rm cloud}/M_{\odot})\sim 4$–6.5. Uncertainty analysis, based on a One-at-a-Time (OAT) methodology, decomposes the total SFR variance into parameter-specific contributions using the coefficient of variation ($\overline{\rm CV}_{i}$). Table.2 shows that the cloud-scale star formation efficiency (equation 11) accounts for $\approx$ 31% of the total uncertainty, due to its weak correlation with cloud mass.
Current limitations of the model include simplifying assumptions regarding clump independence, the adoption of a single-slope CLMF, uniform distributions for key parameters, and a fixed starting time for the embedded phase. Observational constraints on total clump and stellar masses—particularly in low-mass clouds—remain critical for refining predictions. Nevertheless, the model provides a coherent framework linking cloud and clump properties to cloud-scale SFRs and helps guide future observations toward parameters that significantly contribute to predictive uncertainty.
5 Data availability
All data used in this work are available from the first author upon request. The implementation code and its documentation are available online  [^2] https://github.com/jianwenzhou11/SFR_model.git.